# A quick probability-oriented introduction to operator splitting methods

### v1

@@ 12. General results for SPDEs and SDEs
We start with recalling the setting of the filtration theory, as many advances in operator splitting methods for SPDEs start here and it can be considered as a motivating example therefore.
@@
A basic version of the nonlinear filtration problem is as follows. Consider the state $x$ and the noisy measurement $y$ given via
@@
We are going to list three fundamental and general results on the Lie-Trotter splitting for SPDEs.
@@
([31]; see also [32] for a direct proof for the Zakai equation; cf. [323, 119]) Consider
@@
where $W$ is a cylindrical Wiener process and $A,B$ satisfy standard assumptions: hemicontinuity, monotonicity, coercivity for the deterministic part, Lipschitz continuity for the stochastic integral, measurability and controlled growth for either etc. To define a splitting scheme, two equations in some Hilbert space $H$
@@
$dy(t)$ $=B(t,x(t))dW(t)$
are considered in the sense of the variational formulation (through the theory of Gelfand triples), and the actions of their propagators are iterated so the deterministic (a PDE) and stochastic (a Markov semigroup) parts of the original equation are combined exactly as in the Lie-Trotter splitting. Define $t_{k}=\tfrac{k}{n},k=\overline{1,n},$ for some fixed natural $n$ and let $d_{*}(t)$ ($d^{*}(t)$) be the smallest (largest) $t_{k}$ such that $t_{k}\leq t$ ($t_{k}>t$). Then consider [^47] Schemes that swap propagators or combine them on the second stage are also considered.
@@
[155] discusses a variety of SPDEs including those with non-classical assumptions and applications of the two step Lie-Trotter splitting to such evolution equations. In contrast to Example 44, also schemes for which both steps contain stochastic integrals (w.r.t. to orthogonal noises) are considered. Splitting schemes are used to derive a constructive proof of the existence of the initial SPDE, in particular. Splitting is also used to study invariant sets and to obtain comparison theorems (and then establish conditions for positivity of solutions) (see also [216] for such results as corollaries of the classical formulation of the Trotter-Kato formula). Applications include some types of McKean-Vlasov equations.
@@
Using splitting to prove existence theorems is a useful trick (see some references in [155]; also [295, 103, 211, 263, 258] at least to some extend).
@@
A drastically different approach based primarily on probabilistic arguments is developed in [165, 164] where general second-order nonlinear SPDEs
@@
under the assumption of stochastic parabolicity are considered. Here $W$ is a finite-dimensional Wiener process, $v$ is coordinate-wise increasing continuous adapted process, $f,g$ are weakly continuous functions with values in some Sobolev space and $A,L$ are second order differential operators with sufficiently smooth coefficients. The part with stochastic integrals being properly separated from the rest of an equation, convergence for the multistep Lie-Trotter splitting is established. Results about the rate of the convergence for $\operatorname{E}\sup_{t}\|u(t)-u_{n}(t)\|_{H},$ where $H$ is some Sobolev space, are given, in particular.
All three examples use the notion of a weak (variational) solution.
Examples 44, 45 and 46 are very versatile and have an extremely wide range of possible applications, but they are not the only works on the topic. Other references that deal with the theory of nonlinear filtration and, in some cases, provides bounds for the rate of convergence include [131, 187, 228, 84, 349, 258, 103, 21, 233]. In particular, [202] studies splitting for the Kushner equation.
Other results in abstract settings without direct invocations of the filtration theory include [277, 82, 211, 30, 120, 215]. [22] considers the Douglas-Rachford splitting in the variational setting and thus gives a rather rare example of additive splitting in this context.
@@
([131, 228]) Consider the following modification of (43):
@@
[131] also contains a discussion of the practical availability of such schemes, which is a rather typical additional problem in practice.
Mild solutions of SPDEs and the Trotter-Kato formulae are studied in [158, 157, 160, 159] (see also references therein).
@@
where $W$ is a cylindrical Wiener process or a $Q-$Wiener process. Here the nonlinear part $u-u^{3}$ only satisfies the one-sided Lipschitz condition and grows faster than linearly at infinity. It should be noted that the one-sided Lipschitz condition is an example of a non-Lipshitz assumption for SDEs that usually guarantees stability of numerical schemes and existence of moments of the solution.
@@
In particular, this method can be used to preserve ergodic properties of the initial equation even for $a\not\in Lip({\mathbb{R}}^{m})$ (see also [59, 78, 53, 2, 91, 3, 254, 5] for results about structure-preservation properties of splitting methods in the stochastic setting, including those about invariant measures and symplectic properties and e.g. [173] for a deterministic result outside the theory of symplectic integrators).
@@
([56, 57, 58]) The following splitting scheme for the stochastic Allen-Cahn equation is proposed:
@@
The splitting scheme of Example 50 is the Euler-Maruyama scheme of some auxiliary equation. Such an idea appears also in [350, 54, 218].
Some other works on the split-step backward Euler method and similar schemes for SPDEs and SDEs are [59, 350, 54, 53, 110, 137, 218, 90, 89, 333, 332, 12, 7, 199, 219, 110, 229]484848not necessarily with problematic coefficients and in a wide range of situations. [137, 89, 334, 55] combine splitting with space discretization.
Some applications, including those to particle systems, finance, Heston and Langevin models and physics, are [78, 76, 2, 91, 3, 142, 51, 254, 298, 239, 232, 231, 242, 74, 4, 209, 34, 97, 322, 109, 257, 132, 304, 1, 295] [^49] This occasionally includes exponential splitting such as the Ninomiya-Victoir scheme..
Operator splitting methods are well known in the rough path theory. We refer to [185] for a general survey and literature (also [148, Chapter 15], [135, 182], both for RDEs and RPDEs). [133] uses a related approach.
Unfortunately, statistical applications of operator splitting methods seem to be mostly undeveloped. See [287, 6] for some results and references.
@@ 13. A few examples: diffeomorphic flows of SDEs, Brownian web and non-homeomorphic Harris flows
The Itô formula cannot be applied to the processes constructed via this scheme, in contrast to, say, the Euler-Maruyama scheme. A method to overcome this difficulty is proposed in [165] where $y_{n}$ and $u_{n}$ are combined to introduce an approximate process that satisfies some SPDE. For that, given $u_{n},y_{n}$ on the uniform partition of $[0;T]$ define a process $z_{n}$ on $[0;2T]$ via
@@
If $a,\sigma_{j},j=\overline{1,m},$ are Lipschitz continuous, the SPDE setting in [31] applies, and the rate of convergence is known:
@@
If we consider the one-dimensional case and Holder continuous coefficients, the version of the Yamada-Watanabe method developed in [166] can be used to estimate the rate of the convergence.
@@
We can use the classical proof of the Talay-Tubaro expansion [162], the only difference being that we need to use pathwise Taylor expansions instead of the Itô formula. E.g. in the one dimensional case, this yields, for smooth $f$ with bounded derivatives,
@@
([108, 107]) An extreme example is the Brownian web $\{B_{s,t}\},$ a system of Wiener processes in ${\mathbb{R}}$ that start from all points of ${\mathbb{R}}$ at all times, move independently until they meet and merge afterward. Here $B_{s,\cdot}(x)$ is a Wiener process started at $x$ at time $s.$ Though the Brownian web does not admit a representation as a SDE, the Lie-Trotter splitting can still be defined on the interval $[\frac{k}{n};\frac{k+1}{n})$ as
@@
([327]) A Harris flow $X_{s,t}$ is a flow on ${\mathbb{R}}$ in which
@@
for some symmetric strictly positive definite $\phi\in C({\mathbb{R}}).$ Under proper assumptions $X_{s,t},t>s,$ is a piecewise constant function. Moreover, though $X_{s,\cdot}(x)$ admits a representation as a solution to some SDE, the process $X_{s,\cdot}(x)$ may not be the unique solution [336]. Still, one defines the Lie-Trotter splitting as in (54) and the resulting flows converge in distribution to the flow with the same $\phi$ and the corresponding drift. Pushforward measures and dual flows in the reversed time are also convergent.
@@ Acknowledgments
This work was supported by a grant from the Simons Foundation (1030291, M.B.V.).

### v2

@@ 12. General results for SPDEs and SDEs
We start with recalling the setting of the filtration theory, as many advances in operator splitting methods for SPDEs start here and it can be considered as a motivating example.
@@
A standard version of the nonlinear filtration problem is as follows. Consider the state $x$ and the noisy measurement $y$ given via
@@
We are going to list three fundamental and general results for the Lie-Trotter splitting for SPDEs.
@@
([BenGloRas92Approximations]; see also [BenGloRas90approximation] for a direct proof for the Zakai equation; cf. [Te68stabilite, EisenHan22variational]) Consider
@@
where $W$ is a cylindrical Wiener process and $A,B$ satisfy standard assumptions: hemicontinuity, monotonicity, coercivity for the deterministic part, Lipschitz continuity for the stochastic part, measurability and controlled growth for either etc. To define a splitting scheme, two equations in some Hilbert space $H$
@@
$dy(t)$ $=B(t,y(t))dW(t)$
are considered in the variational formulation (via the theory of Gelfand triples), and the actions of their propagators are iterated so the deterministic (a PDE) and stochastic (a Markov semigroup) parts of the original equation are combined exactly as in the Lie-Trotter splitting. Define $t_{k}=\tfrac{k}{n},k=\overline{1,n},$ for some fixed natural $n$ and let $d_{*}(t)$ ($d^{*}(t)$) be the smallest (largest) $t_{k}$ such that $t_{k}\leq t$ ($t_{k}>t$). Then consider [^54] Schemes that swap propagators or combine them on the second stage are also considered.
@@
[GonKo98Fractional] discusses a variety of SPDEs including those with non-classical assumptions and applications of the two step Lie-Trotter splitting to such evolution equations. In contrast to Example 45, also schemes for which both steps contain stochastic integrals (w.r.t. to orthogonal noises) are considered. Splitting schemes are used to derive a constructive proof of the existence of the initial SPDE, in particular. Splitting is also used to study invariant sets and to obtain comparison theorems (and then establish conditions for positivity of solutions) (see also [Ko92comparison] for such results as corollaries of the classical formulation of the Trotter-Kato formula). Applications include some types of McKean-Vlasov equations.
@@
Using splitting to prove existence theorems is a useful trick (see some references in [GonKo98Fractional]; also [San13splitting_up, DeuSan14convergence, KoNo16time, NeidSteZa20convergence_2, Na95remarks] at least to some extend).
@@
A different approach based primarily on probabilistic arguments is developed in [GyKry03Splitting, GyKry03rate] where general second-order nonlinear SPDEs
@@
under the assumption of stochastic parabolicity are considered. Here $W$ is a finite-dimensional Wiener process, $v$ is coordinate-wise increasing continuous adapted process, $f,g$ are weakly continuous functions with values in some Sobolev space and $A,L$ are second order and first order respectively differential operators with sufficiently smooth coefficients. With stochastic integrals being properly separated from the rest of the equation, convergence for the multistep Lie-Trotter splitting is established. Results about the rate of the convergence for $\operatorname{E}\sup_{t}\|u(t)-u_{n}(t)\|_{H},$ where $H$ is some Sobolev space, are given, in particular.
All three examples use the notion of the weak solution.
Examples 45, 46 and 47 are very versatile and cover an extremely wide range of applications, but they are not the only works on the topic. Other references that deal with the theory of nonlinear filtration and, in some cases, provide bounds for the rate of convergence include [FlorLeGland91time_discretization, HopWong86Lie_Trotter, LeGland92splitting_up, CriLobbeSal22application, ZhangZouZhang22splitting_up, Na95remarks, DeuSan14convergence, BagAnLar23energy, LiWangYauZhang23solving]. In particular, [ItoRo00approximation] studies splitting for the Kushner equation.
Other results in an abstract setting without direct invocation of the filtration theory include [PadSheng19convergence, CoxNeer10convergence, KoNo16time, BeMa23convergenceArx, EisenStill22randomizedArx, KofLeMeOs21splitting]. [BarRock17splitting] considers the Douglas-Rachford splitting in the variational setting and thus gives a rather rare example of additive splitting in this context.
@@
([FlorLeGland91time_discretization, LeGland92splitting_up]) Consider the following modification of (44):
@@
[FlorLeGland91time_discretization] also contains a discussion of the practical availability of such schemes, which is a rather typical additional problem in practice.
Mild solutions of SPDEs and the Trotter-Kato formulae are studied in [Go06Trotter, Go15Trotter, Go20weak, Go21Trotter] (see also references therein).
@@
where $W$ is a cylindrical Wiener process or a $Q-$Wiener process. Here the nonlinear part $u-u^{3}$ only satisfies the one-sided Lipschitz condition and grows faster than linearly at infinity [^55] The one-sided Lipschitz condition is an example of a non-Lipshitz assumption for SDEs that usually guarantees stability of numerical schemes and existence of moments of the solution..
@@
In particular, this method can be used to preserve ergodic properties of the initial equation even for $a\not\in Lip({\mathbb{R}}^{m})$ (see also [BuckSamTam22Splitting, CheMelTu21diffusion, BreCo22strong, AbVilZy15long, CuiHongSun21structureArx, AbBuck16splitting, MilTre03quasi, AgaMaMe23random] for results about structure-preservation properties of splitting methods in the stochastic setting, including those about invariant measures and symplectic properties and e.g. [HanKraOs12second] for a deterministic result outside the theory of symplectic integrators).
@@
([BreCuiHong19strong, BreGou19analysis, BreGou20weak]) The following splitting scheme for the stochastic Allen-Cahn equation on the interval $[0;T]$ is proposed:
@@
The splitting scheme of Example 51 is the Euler-Maruyama scheme of some auxiliary equation. Such an idea appears also in [ZhaoJu08convergence, BreCoGior24splitting, KoLarLind18discretization].
Some other works on the split-step backward Euler method and similar schemes for SPDEs and SDEs are [BuckSamTam22Splitting, ZhaoJu08convergence, BreCoGior24splitting, BreCo22strong, DoerTeich10semigroup, FuKoLarLind18strong, KoLarLind18discretization, CuiHongLiuZhou19strong, CuiHong19strong_weak, WangLuLiu08two_stage, WangLi10split_step, AnCriOtto23uniformArx, GerJourcCle16Ninomiya, IguYa21operator, KuhNeer04Lie, DoerTeich10semigroup, Lejay18Girsanov],565656not necessarily with problematic coefficients and in a wide range of situations and [FuKoLarLind18strong, CuiHong19strong_weak, WangMeiLiBu19split_step, BreCoUlan23analysisArx] combine splitting with space discretization.
Some applications, including those to particle systems, Heston and Langevin models and physics and finance in general, are [CheMelTu21diffusion, ChenReis22flexible, AbVilZy15long, CuiHongSun21structureArx, AbBuck16splitting, Gei17iterative, Bou17CayleyArx, MilTre03quasi, SbaiJour13high, LordWang22convergenceArx, LenkMa15second_order, LenkMa15weak, MaMon22weak, ChenMar21convergenceArx, AdamsDuReis22operator, KeLord23adaptive, BerDobMon23piecewiseArx, DesDuDuLau14analysis, TeLi20strongArx, DoHan14high, NadZa13approximation, BasTah12strong, Shar03splitting, HoutToi16application, San13splitting_up] [^57] This occasionally includes exponential splitting such as the Ninomiya-Victoir scheme..
Operator splitting methods are well known in the rough path theory. We refer to [HolKarlKenLie10Splitting] for a general survey and literature (also [GlowOshYin17Splitting, Chapter 15], [FrizOber11splitting, He23convergence], both for RDEs and RPDEs). [FosReisStrange23highArx] uses a related approach.
Unfortunately, statistical applications of operator splitting methods seem to be mostly undeveloped. See [PiSamDit22parameterArx, GerJourCle18asymptotics] for some results and references.
@@ 13. A few examples: diffeomorphic flows of SDEs, Brownian web and non-homeomorphic Harris flows
The Itô formula cannot be applied to the processes constructed via this scheme, in contrast to, say, the Euler-Maruyama scheme. A method to overcome this difficulty is proposed in [GyKry03Splitting] where $y_{n}$ and $u_{n}$ are combined to introduce an approximate process that satisfies some SPDE. For that, given $u_{n},y_{n}$ on the uniform partition of $[0;T]$ define a process $z_{n}$ on $[0;2T]$ via
@@
If $a,\sigma_{j},j=\overline{1,m},$ are Lipschitz continuous, the SPDE setting in [BenGloRas92Approximations] applies, and the rate of convergence is known:
@@
If we consider the one-dimensional case and Holder continuous coefficients, the version of the Yamada-Watanabe method developed in [GyRa11Note] can be used to estimate the rate of the convergence.
@@
We can use the classical proof of the Talay-Tubaro expansion [GraTa13stochastic], the only difference being that we need to use pathwise Taylor expansions instead of the Itô formula. E.g. in the one dimensional case, this yields, for smooth $f$ with bounded derivatives,
@@
([DoVov18Arratia, DorVov20ApproximationsEng]) An extreme example is the Brownian web $\{B_{s,t}\},$ a system of Wiener processes in ${\mathbb{R}}$ that start from all points of ${\mathbb{R}}$ at all times, move independently until they meet and merge afterward. Here $B_{s,\cdot}(x)$ is a Wiener process started at $x$ at time $s.$ Though the Brownian web does not admit a representation as a SDE, the Lie-Trotter splitting can still be defined on the interval $[\frac{k}{n};\frac{k+1}{n})$ as
@@
([Vo22SplittingManu]) A Harris flow $X_{s,t}$ is a flow of one-dimensional correlated Wiener processes in which
@@
for some symmetric strictly positive definite $\phi\in C({\mathbb{R}}).$ Under proper assumptions $X_{s,t},t>s,$ is a piecewise constant function. Moreover, though $X_{s,\cdot}(x)$ admits a representation as a solution to some SDE, the process $X_{s,\cdot}(x)$ may not be the unique solution [WarWa04Spectra]. Still, one defines the Lie-Trotter splitting as in (55) and the resulting flows converge in distribution to the flow with the same $\phi$ and the corresponding drift [^59] Proper numerical schemes for such non-homeomorphic flows are studied in [Gli12Disordering, Nish11Discrete].. Pushforward measures and dual flows in the reversed time are also convergent.
@@ Acknowledgments
[…]